OpenStudy (anonymous):

Simplify (sin^2(x))/(1+(sin^2(x))+(sin^4(x))+(sin^6(x))+(sin^8(x))+...)

OpenStudy (anonymous):

\[ \Large \sum_{i=1}^n\left(\sin^2(x)\right)^{n-1}\] Seems to be geometric: \[\Large s_n=\frac{1-q^n}{1-q}=\frac{1-\sin^{2n}(x)}{1-\sin^2(x)}=\frac{1-\sin^{2n}(x)}{\cos^2(x)} \]

OpenStudy (anonymous):

for the denominator of your given term.

OpenStudy (anonymous):

The rest I would guess is a matter of art and interpretation, for large numbers of n, your sin^(2n)(x) cycles between 0 and 1, so it is possible for that term to converge, if that's the case your answer would be: \[\Large \sin^2(x)\cos^2(x) \]

OpenStudy (anonymous):

So written out it might be more informative: \[\Large \sum_{n=1}^\infty\left(\sin^2(x)\right)^{n-1}=\frac{1}{\cos^2(x)}=\sec^2(x) \] And \[\Large \frac{\sin^2(x)}{\sum_{n=1}^\infty\left(\sin^2(x)\right)^{n-1}}=\frac{\sin^2(x)}{\sec^2(x)}=\sin^2(x)\cos^2(x) \]

OpenStudy (anonymous):

Thanks! To add on, Since \[2\sin(x)\cos(x)=\sin(2x)\] Then \[\sin^2(x)\cos^2(x)=(\sin(x)\cos(x))^2=(1/2\sin(2x))^2=1/4\sin^2(2x)\]

OpenStudy (anonymous):

Yes, very true! Makes it even more elegant. (-: \[ \Huge \checkmark\]